OpenStudy (anonymous):

what would be the % of 32 out of 50??

OpenStudy (hba):

32/2

mathslover (mathslover):

ok ! so .. let the required % be x

mathslover (mathslover):

x% of 32 = 50 \(\frac{x}{100}*32=50\)

mathslover (mathslover):

can u solve for x now @homeisawsome ?

OpenStudy (anonymous):

@mathslover u are very Quick in typing equation!

mathslover (mathslover):

thanks @sami-21 ... @homeisawsome did u get it ?? : \[\frac{x}{100}*32=50\] \[\frac{x}{\cancel{100}50}*\cancel{32}16=50\] \[\frac{x}{\cancel{50}25}*\cancel{16}8=50\]

OpenStudy (anonymous):

What is your answer @homeisawsome ??

mathslover (mathslover):

yes @homeisawsome let us know about that .. if u dont know the answer .. let us know that where r u having problem ?

OpenStudy (anonymous):

32%-50 ?

OpenStudy (anonymous):

@homeisawsome are you listening or not??

OpenStudy (anonymous):

just a beginner.

OpenStudy (anonymous):

thanks and sorry my comp froze

mathslover (mathslover):

no problem @homeisawsome \[\Huge{\textbf{Best}\mathbb{Of}\textbf{Luck}}\]

OpenStudy (anonymous):

what is the right approach to the question in hand

OpenStudy (anonymous):

can anybody please help